OpenStudy (anonymous):

The height of an equilateral triangular prism increases by four units. The new volume is more than the original by how much? four times the area of the base four more than the the area of the base four more than the length of the height four times the length of the height

OpenStudy (anonymous):

a

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

wat?

OpenStudy (anonymous):

nothing

OpenStudy (anonymous):

k

OpenStudy (cwrw238):

volume of the prism = height * area of base = h*a so if base is 4 units more new volume = (h+4) * a

OpenStudy (cwrw238):

= ha + 4a

OpenStudy (anonymous):

so its not wat she said?

OpenStudy (cwrw238):

its 4 times area of the base

OpenStudy (cwrw238):

yes thats what cemantha said

OpenStudy (anonymous):

ok thanks alot